Analysis of Variance Table Response: Price

Df Sum Sq Mean Sq F value Pr(>F)

Living.Area 1 1.3501e+12 1.3501e+12 362.0394 < 2e-16 ***

Bedrooms 1 2.3642e+10 2.3642e+10 6.3394 0.01241 *

Fireplaces 1 7.6232e+07 7.6232e+07 0.0204 0.88642

Residuals 259 9.6588e+11 3.7293e+09

--- Signif. codes: 0 ‘***’ 0.001 ‘**’ 0.01 ‘*’ 0.05 ‘.’ 0.1 ‘ ’ 1 >

Using= 0.05, perform an F test of overall linear relationship. State the hypotheses, the value of F-test statistic, p-value, and your conclusion.
Answer #1

HI : βί t0for atleast one i

The test is based on the statistic F=[SS(Regression)/3]/[SS(Residual)/259]

Then under the null, F~F3,259

From the output,

SS(Regression)=1.3501e+12+2.3642e+10+7.6232e+07=1.373818e+12
SS(Residual)=9.6588e+11

Observed F=(1.373818e+12/3)/(9.6588e+11/259) =122.7961

F.05,3,259=2.639455

p value=P(F3,259>122.7961)= 0.00000 approximately

Thus the predictor variables under consideration are significantly worthwhile in predicting Price.